# Statistical Test on Diffusion Model-based Generated Images by Selective Inference

### v2

@@ Abstract
AI technology for generating images, such as diffusion models, has advanced rapidly. However, there is no established framework for quantifying the reliability of AI-generated images, which hinders their use in critical decision-making tasks, such as medical image diagnosis. In this study, we propose a method to quantify the reliability of decision-making tasks that rely on images produced by diffusion models within a statistical testing framework. The core concept of our statistical test involves using a selective inference framework, in which the statistical test is conducted under the condition that the images are produced by a trained diffusion model. As a case study, we study a diffusion model-based anomaly detection task for medical images. With our approach, the statistical significance of medical image diagnostic outcomes can be quantified in terms of a $p$-value, enabling decision-making with a controlled error rate. We demonstrate the theoretical soundness and practical effectiveness of our statistical test through numerical experiments on both synthetic and brain image datasets.
@@ 1 Introduction
On the other hand, when using virtual images generated by AI for critical decision-making, such as medical diagnosis, it is crucial to ensure the reliability of the decisions. Given that images are generated by an AI algorithm, such as a deep learning model trained on historical data, it may inherently contain biases and errors. Therefore, treating virtual synthetic images as equivalent to real images in decision-making tasks carries the risk of biased and erroneous outcomes. When making critical decisions based on generated images, it is necessary to be able to assess their reliability by properly taking into account the fact that the images were generated by AI. However, to our knowledge, there are no studies that can quantify the reliability of decision-making based on image generation AI.
Figure 1: Schematic illustration of the anomaly detection on a brain image dataset using a diffusion model and the proposed DMAD-test. When a test image, which may contain an abnormal region, is fed into a trained diffusion model, a normal image is generated through the forward process and reverse process. By initiating the image generation from the middle of the forward process, a normal image that retains the characteristics of the input image can be generated. By comparing the input image with the normal image, the anomaly region can be identified. In this study, we propose a method called the DMAD-test, which quantifies the statistical significance of the identified anomaly regions in the form of $p$-value. The proposed DMAD-test calculates the $p$-values by incorporating the fact that the anomaly region has been identified by the diffusion model, thus enabling unbiased decision-making (see §3 and §4 for details).
@@
As an example of decision-making tasks based on image generation AI, we focus on the problem of detecting anomaly regions in medical images (Wolleb et al., 2022; Baur et al., 2021) (see Figure 1). Initially, a diffusion model is trained exclusively on normal images during the training phase. In the testing phase, a patient’s test image is processed through this model to create a virtual normal image, against which the original is compared to identify anomaly regions. Our proposed statistical test, the Diffusion Model-based Anomaly Region Detection Test (DMAD-test), quantifies the statistical reliability of detected anomalies as $p$-values. Decisions based on these $p$-values can theoretically control the false detection rate at desired significance levels (such as 0.01 or 0.05) even in finite sample settings.
Related works.
@@
Diffusion models have been effectively utilized in anomaly region detection problems (Wolleb et al., 2022; Pinaya et al., 2022; Fontanella et al., 2023; Wyatt et al., 2022; Mousakhan et al., 2023). In this context, the denoising diffusion probabilistic model (DDPM) is commonly used (Ho et al., 2020; Song et al., 2022). During the training phase, a DDPM model learns the distribution of normal medical images by iteratively adding and then removing noise. In the test phase, the model attempts to reconstruct a new test image. If the image contains abnormal regions, such as tumors, the model may struggle to accurately reconstruct these regions, as it has been trained primarily on normal regions. The discrepancies between the original and the reconstructed image are then analyzed to identify and highlight anomaly regions. Other types of generative AI has also been used for anomaly region detection task (Baur et al., 2021; Chen and Konukoglu, 2018; Chow et al., 2020; Jana et al., 2022).
@@
SI was first introduced within the context of reliability evaluation for linear model features when they were selected using a feature selection algorithm (Lee and Taylor, 2014; Lee et al., 2016; Tibshirani et al., 2016), and then extended to more complex feature selection methods (Yang et al., 2016; Suzumura et al., 2017; Hyun et al., 2018; Rügamer and Greven, 2020; Das et al., 2021). Then, SI proves valuable not only for feature selection problems but also for statistical inference across various data-driven hypotheses, including unsupervised learning tasks (Chen and Bien, 2020; Tsukurimichi et al., 2021; Tanizaki et al., 2020; Duy et al., 2022; Le Duy et al., 2024; Lee et al., 2015; Gao et al., 2022; Duy et al., 2020; Jewell et al., 2022). The fundamental idea of SI is to perform an inference conditional on the hypothesis selection event, which mitigates the selection bias issue even when the hypothesis is selected and tested using the same data. To conduct SI, it is necessary to derive the sampling distribution of test statistic conditional on the hypothesis selection event. To the best of our knowledge, SI was applied to statistical inferences on several deep learning models (Duy et al., 2022; Miwa et al., 2023; Shiraishi et al., 2024; Miwa et al., 2024), but none of them work on image generation by diffusion models.
@@ 2 Diffusion Models
In this section, we briefly explain the diffusion model employed in this study. Given a test image which possibly contain abnormal regions, a denoising diffusion model (Ho et al., 2020; Song et al., 2022) is used to generate the corresponding normal image. The reconstruction process consists of two processes called forward process (or diffusion process) and reverse process.
In the forward process, noise is sequentially added to the test image so that it converges to a standard Gaussian distribution $\mathcal{N}(\bm{0},\bm{I})$. Let $\mathbf{x}$ be an image represented as a vector with each element corresponding to a pixel value. Given an original test image $\mathbf{x}_{0}$, noisy images $\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{T}$ are sequentially generated, where $T$ is the number of noise addition steps. We consider the distribution of the original and noisy test images, which is denoted by $q(\mathbf{x})$, and approximate the distribution by a parametric model $p_{\theta}(\mathbf{x})$ with $\theta$ being the parameters. Using a sequence of noise scheduling parameters $0<\beta_{1}<\beta_{2},<\dots<\beta_{T}<1$, the forward process is written as
$q(\mathbf{x}_{1:T}|\mathbf{x}_{0}):=\prod_{(t\geq 1)}^{T}q(\mathbf{x}_{t}|\mathbf{x}_{t-1}),$
where
$q(\mathbf{x}_{t}|\mathbf{x}_{t-1}):=\mathcal{N}(\mathbf{x}_{t};\sqrt{1-\beta_{t}}\mathbf{x}_{t-1},\beta_{t}\bm{I}).$
@@
$\mathbf{x}_{t}=\sqrt{\alpha_{t}}\mathbf{x}_{0}+\sqrt{1-\alpha_{t}}\epsilon,\;\;\;\text{with}\;\;\;\epsilon\sim\mathcal{N}(\mathbf{0},\bm{I}),$ (1)
@@
In the reverse process, a parametric model in the form of $p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})=\mathcal{N}(\mathbf{x}_{t-1};\mu_{\theta}(\mathbf{x}_{t},t),\beta_{t}\bm{I})$ is employed, where $\mu_{\theta}(\mathbf{x}_{t},t)$ is obtained by using the predicted noise component $\epsilon^{(t)}_{\theta}(\mathbf{x}_{t})$. Typically, a U-Net is used as the model architecture for $\epsilon^{(t)}_{\theta}(\mathbf{x}_{t})$. In DDPM (Ho et al., 2020), the loss function for training the noise component is simply written as $||\epsilon^{(t)}_{\theta}(\mathbf{x}_{t})-\epsilon||^{2}_{2}$. Based on (1), given a noisy image $\mathbf{x}_{t}$ after $t$ steps, the reconstruction of the image in the previous step $\mathbf{x}_{t-1}$ is obtained as
@@
$\sigma_{t}=\eta\sqrt{\frac{1-\alpha_{t-1}}{1-\alpha_{t}}}\sqrt{1-\frac{\alpha_{t}}{\alpha_{t-1}}}.$ (4)
@@
In practice, the reverse process starts from $\mathbf{x}_{T^{\prime}}$ with $T^{\prime}<T$. Namely, we reconstruct the original input image not from the completely noisy one, but from a one which still contain individual information of the original input image. The smaller $T^{\prime}$ ensures that the reconstructed image preserves fine details of the input image. Conversely, the larger $T^{\prime}$ results in the retention of only large scale features, thereby converting more of the abnormal regions into normal regions (Ho et al., 2020; Mousakhan et al., 2023). Therefore, $T^{\prime}$ should be set to balance the feature retention of the input image and the conversion of the abnormal region to the normal region. Note that setting $T^{\prime}$ smaller than $T$ has advantages in terms of computational cost. For the purpose of reducing computational cost, various methods have been proposed. For example, one way is to sample while skipping portions of the sampling trajectory (see Appendix A). The image reconstruction scheme by DDPM is summarized in Algorithm 1.
@@
1:  $\mathbf{x}_{T^{\prime}}\leftarrow\sqrt{\alpha_{T^{\prime}}}\mathbf{x}+\sqrt{1-\alpha_{T^{\prime}}}\epsilon_{T^{\prime}}$
@@ 3 Statistical Test on Generated Images by Diffusion Models
In this section, we formulate the statistical test for detecting anomaly regions using images generated by a trained DDPM model. As shown in Figure 1, anomaly region detection by diffusion models is conducted as follows. First, in the training phase, the diffusion model is trained only on normal images. Then, in the test phase, we feed a test image which might contain anomaly regions into the trained diffusion model, and reconstruct it back from a noisy image $\mathbf{x}_{T^{\prime}}$ at step $T^{\prime}<T$. By appropriately selecting $T^{\prime}$, we can generate a normal image that retain individual characteristics of the test input image. If the image does not contain anomaly regions, the reconstructed image is expected to be similar to the original test image. On the other hand, if the image contains anomaly regions, such as tumors, the model may struggle to accurately reconstruct these regions, as it has been trained primarily on normal regions. Therefore, the anomaly regions can be detected by comparing the original test image and its reconstructed one.
@@
We develop a statistical test to quantify the reliability of decision-making based on images generated by diffusion models. To develop a statistical test, we interpret that an image is considered as a sum of true signal component $\bm{\mu}\in\mathbb{R}^{n}$ and noise component $\bm{\epsilon}\in\mathbb{R}^{n}$. Regarding the true signal component, each pixel can have an arbitrary value without any particular assumption or constraint. On the other hand, regarding the noise component, it is assumed to follow a Gaussian distribution, and their covariance matrix is estimated using normal data different from that used for the training of the diffusion model, which is the standard setting of SI. Namely, an image with $n$ pixels can be represented as an $n$-dimensional random vector
$\bm{X}=(X_{1},X_{2},\dots,X_{n})^{\top}=\bm{\mu}+\bm{\epsilon},\;\;\;\bm{\epsilon}\sim\mathcal{N}(\bm{0},\Sigma),$ (5)
where $\bm{\mu}\in\mathbb{R}^{n}$ is the unknown true signal vector and $\Sigma$ is the covariance matrix, which is known or estimable from external data. In the following, we use capital $\bm{X}$ to emphasize that an image is considered as a random vector, while the observed image is denoted as $\bm{X}^{\rm obs}$.
Let us denote the reconstruction process of the trained diffusion model in Algorithm 1 as the mapping from an input image to the reconstructed image $\mathcal{D}:\mathbb{R}^{n}\ni\bm{X}\rightarrow\mathcal{D}(\bm{X})\in\mathbb{R}^{n}$. The difference between the input image $\bm{X}$ and the reconstructed image $\mathcal{D}(\bm{X})$ indicates the reconstruction error. When identifying anomaly regions based on reconstruction error, it is useful to apply some filter to remove the influence of pixel-wise noise. In this study, we simply used an averaging filter. Let us denote the averaging filter as $\mathcal{F}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$. Then, the process of obtaining the (filtered) reconstruction error is written as
@@
where absolute value is taken pixel-wise. Anomaly regions are then obtained by applying a threshold to the filtered reconstruction error $E_{i}(\bm{X})$ for each pixel $i\in[n]$. Specifically, we define the anomaly region as the set of pixels whose filtered reconstruction error is greater than a given threshold $\lambda\in(0,\infty)$, i.e.,
@@
In order to quantify the statistical significance of the anomaly regions detected by a diffusion model, let us consider a statistical test with the following null and alternative hypotheses:
$\mathrm{H}_{0}\colon\frac{1}{|\mathcal{M}_{\bm{X}}|}\sum_{i\in\mathcal{M}_{\bm{X}}}\mu_{i}=\frac{1}{|\mathcal{M}_{\bm{X}}^{c}|}\sum_{i\notin\mathcal{M}_{\bm{X}}}\mu_{i}~{}~{}~{}\text{v.s.}~{}~{}~{}\mathrm{H}_{1}\colon\frac{1}{|\mathcal{M}_{\bm{X}}|}\sum_{i\in\mathcal{M}_{\bm{X}}}\mu_{i}\neq\frac{1}{|\mathcal{M}_{\bm{X}}^{c}|}\sum_{i\notin\mathcal{M}_{\bm{X}}}\mu_{i},$ (8)
where $\mathrm{H}_{0}$ is the null hypothesis that the mean pixel values are the same between the anomaly regions and other regions, while $\mathrm{H}_{1}$ is the alternative hypothesis that they are different. A reasonable test statistic for the statistical test in (8) is the difference in mean pixel values between the inside and outside of the anomaly region, i.e.,
$T(\bm{X})=\frac{1}{|\mathcal{M}_{\bm{X}}|}\sum_{i\in\mathcal{M}_{\bm{X}}}X_{i}-\frac{1}{|\mathcal{M}^{c}_{\bm{X}}|}\sum_{i\in\mathcal{M}^{c}_{\bm{X}}}X_{i}=\bm{\nu}^{\top}_{\mathcal{M}_{\bm{X}}}\bm{X},$ (9)
where $\bm{\nu}_{\mathcal{M}_{\bm{X}}}=\frac{1}{|\mathcal{M}_{\bm{X}}|}\bm{1}^{n}_{\mathcal{M}_{\bm{X}}}-\frac{1}{|\mathcal{M}^{c}_{\bm{X}}|}\bm{1}^{n}_{\mathcal{M}^{c}_{\bm{X}}}$ is the vector that depends on the anomaly region $\mathcal{M}_{\bm{X}}$ and $\mathbf{1}^{n}_{\mathcal{C}}\in\mathcal{R}^{n}$ is an $n$-dimensional vector whose elements are $1$ if they belong to the set $\mathcal{C}$ and 00 otherwise.
If we do not account for the fact that the anomaly region is detected by a diffusion model, the distribution of the test statistic would be simply given as $T(\bm{X})\sim\mathcal{N}(0,\bm{\nu}^{\top}_{\mathcal{M}_{\bm{X}}}\Sigma\bm{\nu}_{\mathcal{M}_{\bm{X}}})$. In this case, the $p$-values defined as
$p_{\mathrm{naive}}=\mathbb{P}_{\mathrm{H}_{0}}\left(|T(\bm{X})|>|T(\bm{X}^{\mathrm{obs}})|\right),$ (10)
would be easily computed by the normality of the test statistic distribution. However, in reality, since the anomaly region is detected by the trained diffusion model, $\bm{\nu}_{\mathcal{M}_{\bm{X}}}$ depends on the data $\bm{X}$, meaning that the sampling distribution of the test statistic is much more complicated. Therefore, if $p_{\mathrm{naive}}$ is used for decision-making, the false detection error cannot be properly controlled.

### v3

# Statistical Test on Diffusion Model-based Anomaly Detection by Selective Inference

@@ Abstract
Advancements in AI image generation, particularly diffusion models, have progressed rapidly. However, the absence of an established framework for quantifying the reliability of AI-generated images hinders their use in critical decision-making tasks, such as medical image diagnosis. In this study, we address the task of detecting anomalous regions in medical images using diffusion models and propose a statistical method to quantify the reliability of the detected anomalies. The core concept of our method involves a selective inference framework, wherein statistical tests are conducted under the condition that the images are produced by a diffusion model. With our approach, the statistical significance of anomaly detection results can be quantified in the form of a $p$-value, enabling decision-making with controlled error rates, as is standard in medical practice. We demonstrate the theoretical soundness and practical effectiveness of our statistical test through numerical experiments on both synthetic and brain image datasets.
@@ 1 Introduction
On the other hand, when using virtual images generated by AI for critical decision-making, such as medical diagnosis, it is crucial to ensure the reliability of the decisions. Given that images are generated by an AI algorithm, such as a deep learning model trained on historical data, they may inherently contain biases and errors. Therefore, treating virtual synthetic images as equivalent to real images in decision-making tasks carries the risk of biased and erroneous outcomes. When making critical decisions based on generated images, it is necessary to be able to assess their reliability by properly taking into account the fact that the images were generated by AI. However, to our knowledge, there are no studies that can quantify the reliability of decision-making based on image generation AI.
Figure 1: Schematic illustration of the anomaly detection on a brain image dataset using a diffusion model and the proposed DMAD-test. When a test image, which may contain an anomalous region, is fed into a trained diffusion model, a normal image is generated through the forward process and reverse process. By initiating the image generation from the middle of the forward process, a normal image that retains the characteristics of the input image can be generated. By comparing the input image with the normal image, the anomalous region can be identified. In this study, we propose a method called the DMAD-test, which quantifies the statistical significance of the identified anomalous regions in the form of $p$-value. The DMAD-test calculates the $p$-values by incorporating the fact that the anomalous region has been identified by the diffusion model, thus enabling unbiased decision-making (see §3 and §4 for details).
@@
As an example of decision-making tasks based on image generation AI, we focus on the problem of detecting anomalous regions in medical images (Wolleb et al., 2022; Baur et al., 2021) (see Figure 1). Initially, a diffusion model is trained exclusively on normal images during the training phase. In the testing phase, a patient’s test image is processed through this model to create a virtual normal image, against which the original is compared to identify anomalous regions. Our proposed statistical test, the Diffusion Model-based Anomalous Region Detection Test (DMAD-test), quantifies the statistical reliability of detected anomalies as $p$-values. Decisions based on these $p$-values can theoretically control the false detection rate at desired significance levels (such as 0.01 or 0.05).
Related work.
@@
Diffusion models have been effectively utilized in anomalous region detection problems (Wolleb et al., 2022; Pinaya et al., 2022; Fontanella et al., 2023; Wyatt et al., 2022; Mousakhan et al., 2023). In this context, the denoising diffusion probabilistic model (DDPM) is commonly used (Ho et al., 2020; Song et al., 2022). During the training phase, a DDPM model learns the distribution of normal medical images by iteratively adding and then removing noise. In the test phase, the model attempts to reconstruct a new test image. If the image contains anomalous regions, such as tumors, the model may struggle to accurately reconstruct these regions, as it has been trained primarily on normal regions. The discrepancies between the original and the reconstructed image are then analyzed to identify and highlight anomalous regions. Other types of generative AI has also been used for anomalous region detection task (Baur et al., 2021; Chen and Konukoglu, 2018; Chow et al., 2020; Jana et al., 2022).
@@
SI was first introduced within the context of reliability evaluation for linear model features when they were selected using a feature selection algorithm (Lee and Taylor, 2014; Lee et al., 2016; Tibshirani et al., 2016), and then extended to more complex feature selection methods (Yang et al., 2016; Suzumura et al., 2017; Hyun et al., 2018; Rügamer and Greven, 2020; Das et al., 2021). Then, SI proves valuable not only for feature selection problems but also for statistical inference across various data-driven hypotheses, including unsupervised learning tasks (Chen and Bien, 2020; Tsukurimichi et al., 2021; Tanizaki et al., 2020; Duy et al., 2022; Le Duy et al., 2024; Lee et al., 2015; Gao et al., 2022; Duy et al., 2020; Jewell et al., 2022). The fundamental idea of SI is to perform an inference conditional on the hypothesis selection event, which mitigates the selection bias issue even when the hypothesis is selected and tested using the same data. To conduct SI, it is necessary to derive the sampling distribution of test statistic conditional on the hypothesis selection event. To the best of our knowledge, SI was applied to statistical inferences on several deep learning models (Duy et al., 2022; Miwa et al., 2023; Shiraishi et al., 2024; Miwa et al., 2024), but none of them works on image generation by diffusion models.
@@ 2 Diffusion Models
In this section, we briefly explain the diffusion model employed in this study. Given a test image which possibly contain anomalous regions, a denoising diffusion model (Ho et al., 2020; Song et al., 2022) is used to generate the corresponding normal image. The reconstruction process consists of two processes called forward process (or diffusion process) and reverse process.
In the forward process, noise is sequentially added to the test image so that it converges to a standard Gaussian distribution $\mathcal{N}(\bm{0},I)$. Let $\mathbf{x}$ be an image represented as a vector with each element corresponding to a pixel value. Given an original test image $\mathbf{x}_{0}$, noisy images $\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{T}$ are sequentially generated, where $T$ is the number of noise addition steps. We consider the distribution of the original and noisy test images, which is denoted by $q(\mathbf{x})$, and approximate the distribution by a parametric model $p_{\theta}(\mathbf{x})$ with $\theta$ being the parameters. Using a sequence of noise scheduling parameters $0<\beta_{1}<\beta_{2},<\dots<\beta_{T}<1$, the forward process is written as
$q(\mathbf{x}_{1:T}|\mathbf{x}_{0}):=\prod_{t=1}^{T}q(\mathbf{x}_{t}|\mathbf{x}_{t-1}),\quad\text{where}\;\;q(\mathbf{x}_{t}|\mathbf{x}_{t-1}):=\mathcal{N}(\sqrt{1-\beta_{t}}\mathbf{x}_{t-1},\beta_{t}I).$
@@
$\mathbf{x}_{t}=\sqrt{\alpha_{t}}\mathbf{x}_{0}+\sqrt{1-\alpha_{t}}\epsilon,\;\;\;\text{with}\;\;\;\epsilon\sim\mathcal{N}(\mathbf{0},I),$ (1)
@@
In the reverse process, a parametric model in the form of $p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})=\mathcal{N}(\mathbf{x}_{t-1};\mu_{\theta}(\mathbf{x}_{t},t),\beta_{t}I)$ is employed, where $\mu_{\theta}(\mathbf{x}_{t},t)$ is obtained by using the predicted noise component $\epsilon^{(t)}_{\theta}(\mathbf{x}_{t})$. Typically, a U-Net is used as the model architecture for $\epsilon^{(t)}_{\theta}(\mathbf{x}_{t})$. In DDPM (Ho et al., 2020), the loss function for training the noise component is simply written as $||\epsilon^{(t)}_{\theta}(\mathbf{x}_{t})-\epsilon_{t}||^{2}_{2}$. Based on (1), given a noisy image $\mathbf{x}_{t}$ after $t$ steps, the reconstruction of the image in the previous step $\mathbf{x}_{t-1}$ is obtained as
@@
$\sigma_{t}=\eta\sqrt{(1-\alpha_{t-1})/(1-\alpha_{t})}\sqrt{1-\alpha_{t}/\alpha_{t-1}}.$ (4)
@@
In practice, the reverse process starts from $\mathbf{x}_{T^{\prime}}$ with $T^{\prime}<T$. Namely, we reconstruct the original input image not from the completely noisy one, but from a one which still contain individual information of the original input image. The smaller $T^{\prime}$ ensures that the reconstructed image preserves fine details of the input image. Conversely, the larger $T^{\prime}$ results in the retention of only large scale features, thereby converting more of the anomalous regions into normal regions (Ho et al., 2020; Mousakhan et al., 2023). Therefore, $T^{\prime}$ should be set to balance the feature retention of the input image and the conversion of the anomalous region to the normal region. Note that setting $T^{\prime}$ smaller than $T$ has advantages in terms of computational cost. For the purpose of reducing computational cost, various methods have been proposed. For example, one way is to sample while skipping portions of the sampling trajectory (see Appendix A). The image reconstruction scheme by DDPM is summarized in Algorithm 1.
@@
1:  $\mathbf{x}_{T^{\prime}}\leftarrow\sqrt{\alpha_{T^{\prime}}}\mathbf{x}+\sqrt{1-\alpha_{T^{\prime}}}\epsilon$
@@ 3 Statistical Test on Generated Images by Diffusion Models
In this section, we formulate the statistical test for detecting anomalous regions using images generated by a trained DDPM model. As shown in Figure 1, anomalous region detection by diffusion models is conducted as follows. First, in the training phase, the diffusion model is trained only on normal images. Then, in the test phase, we feed a test image which might contain anomalous regions into the trained diffusion model, and reconstruct it back from a noisy image $\mathbf{x}_{T^{\prime}}$ at step $T^{\prime}<T$. By appropriately selecting $T^{\prime}$, we can generate a normal image that retain individual characteristics of the test input image. If the image does not contain anomalous regions, the reconstructed image is expected to be similar to the original test image. On the other hand, if the image contains anomalous regions, such as tumors, the model may struggle to accurately reconstruct these regions, as it has been trained primarily on normal regions. Therefore, the anomalous regions can be detected by comparing the original test image and its reconstructed one.
@@
We develop a statistical test to quantify the reliability of decision-making based on images generated by diffusion models. To develop a statistical test, we interpret an image as a sum of a true signal component $\bm{\mu}\in\mathbb{R}^{n}$ and a noise component $\bm{\varepsilon}\in\mathbb{R}^{n}$. We emphasize that the noise component $\bm{\varepsilon}$ should not be confused with the noise $\epsilon$ used in the forward process. Regarding the true signal component, each pixel can have an arbitrary value without any particular assumption or constraint. On the other hand, regarding the noise component, it is assumed to follow a Gaussian distribution, and their covariance matrix is estimated using normal data different from that used for the training of the diffusion model, which is the standard setting of SI. Namely, an image with $n$ pixels can be represented as an $n$-dimensional random vector
$\bm{X}=(X_{1},X_{2},\dots,X_{n})^{\top}=\bm{\mu}+\bm{\varepsilon},\;\;\;\bm{\varepsilon}\sim\mathcal{N}(\bm{0},\Sigma),$ (5)
where $\bm{\mu}\in\mathbb{R}^{n}$ is the unknown true signal vector and $\Sigma$ is the covariance matrix. In the following, we use capital $\bm{X}$ to emphasize that an image is considered as a random vector, while the observed image is denoted as $\bm{X}_{\rm obs}$.
Let us denote the reconstruction process of the trained diffusion model in Algorithm 1 as the mapping from an input image to the reconstructed image $\mathcal{D}:\mathbb{R}^{n}\ni\bm{X}\rightarrow\mathcal{D}(\bm{X})\in\mathbb{R}^{n}$. The difference between the input image $\bm{X}$ and the reconstructed image $\mathcal{D}(\bm{X})$ indicates the reconstruction error. When identifying anomalous regions based on reconstruction error, it is useful to apply some filter to remove the influence of pixel-wise noise. In this study, we simply used an averaging filter. Let us denote the averaging filter as $\mathcal{F}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$. Then, the process of obtaining the (filtered) reconstruction error is written as
@@
where absolute value is taken pixel-wise. Anomalous regions are then obtained by applying a threshold to the filtered reconstruction error $E_{i}(\bm{X})$ for each pixel $i\in[n]$. Specifically, we define the anomalous region as the set of pixels whose filtered reconstruction error is greater than a given threshold $\lambda\in(0,\infty)$, i.e.,
@@
In order to quantify the statistical significance of the anomalous regions detected by a diffusion model, we consider the concrete example of two-sample test. Let us define an $n$-dimensional reference input vector,
$\bm{X}^{\mathrm{ref}}=(X^{\mathrm{ref}}_{1},X^{\mathrm{ref}}_{2},\dots,X^{\mathrm{ref}}_{n})^{\top}=\bm{\mu}^{\mathrm{ref}}+\bm{\varepsilon}^{\mathrm{ref}},\;\;\;\bm{\varepsilon}^{\mathrm{ref}}\sim\mathcal{N}(\bm{0},\Sigma),$ (8)
where $\bm{\mu}^{\mathrm{ref}}\in\mathbb{R}^{n}$ is the unknown true signal vector of the reference image and the $\bm{\varepsilon}^{\mathrm{ref}}\in\mathbb{R}^{n}$ is the noise component. Then, we consider the following null and alternative hypotheses:
$\begin{gathered}\mathrm{H}_{0}\colon\frac{1}{|\mathcal{M}_{\bm{X}}|}\sum_{i\in\mathcal{M}_{\bm{X}}}\mu_{i}=\frac{1}{|\mathcal{M}_{\bm{X}}|}\sum_{i\in\mathcal{M}_{\bm{X}}}\mu^{\rm ref}_{i}\\ \text{v.s.}\\ \mathrm{H}_{1}\colon\frac{1}{|\mathcal{M}_{\bm{X}}|}\sum_{i\in\mathcal{M}_{\bm{X}}}\mu_{i}\neq\frac{1}{|\mathcal{M}_{\bm{X}}|}\sum_{i\in\mathcal{M}_{\bm{X}}}\mu^{\rm ref}_{i},\end{gathered}$ (9)
where $\mathrm{H}_{0}$ is the null hypothesis that the mean pixel values are the same between the test image and the reference images in the anomalous regions, while $\mathrm{H}_{1}$ is the alternative hypothesis that they are different. A reasonable test statistic for the statistical test in (9) is the difference in mean pixel values between the test image and the reference image in the anomalous region $\mathcal{M}_{\bm{X}}$, i.e.,
$T(\bm{X},\bm{X}^{\mathrm{ref}})=\frac{1}{|\mathcal{M}_{\bm{X}}|}\sum_{i\in\mathcal{M}_{\bm{X}}}X_{i}-\frac{1}{|\mathcal{M}_{\bm{X}}|}\sum_{i\in\mathcal{M}_{\bm{X}}}X^{\mathrm{ref}}_{i}=\bm{\nu}^{\top}_{\mathcal{M}_{\bm{X}}}\begin{pmatrix}\bm{X}\\ \bm{X}^{\mathrm{ref}}\end{pmatrix},$ (10)
where $\bm{\nu}_{\mathcal{M}_{\bm{X}}}\in\mathbb{R}^{2n}$ is the vector that depends on the anomalous region $\mathcal{M}_{\bm{X}}$, defined as
$\bm{\nu}_{\mathcal{M}_{\bm{X}}}=\frac{1}{|\mathcal{M}_{\bm{X}}|}\begin{pmatrix}\bm{1}^{n}_{\mathcal{M}_{\bm{X}}}\\[5.0pt] -\bm{1}^{n}_{\mathcal{M}_{\bm{X}}}\end{pmatrix}\in\mathbb{R}^{2n},$ (11)
where $\bm{1}^{n}_{\mathcal{C}}\in\mathbb{R}^{n}$ is an $n$-dimensional vector whose elements are $1$ if they belong to the set $\mathcal{C}$ and 00 otherwise. If we do not account for the fact that the anomalous region is detected by a diffusion model, the distribution of the test statistic would be simply given as
$T(\bm{X},\bm{X}^{\mathrm{ref}})\sim\mathcal{N}(0,\bm{\nu}^{\top}_{\mathcal{M}_{\bm{X}}}\tilde{\Sigma}\bm{\nu}_{\mathcal{M}_{\bm{X}}}),\;\;\;\text{where}\;\;\tilde{\Sigma}=\begin{pmatrix}\Sigma&O_{n}\\ O_{n}&\Sigma\end{pmatrix}.$ (12)
In this case, the $p$-values defined as
$p_{\mathrm{naive}}=\mathbb{P}_{\mathrm{H}_{0}}\left(|T(\bm{X},\bm{X}^{\mathrm{ref}})|>|T(\bm{X}_{\mathrm{obs}},\bm{X}^{\mathrm{ref}}_{\mathrm{obs}})|\right),$ (13)
would be easily computed by the normality of the test statistic distribution. However, in reality, since the anomalous region is detected by the trained diffusion model, $\bm{\nu}_{\mathcal{M}_{\bm{X}}}$ depends on the data $\bm{X}$, meaning that the sampling distribution of the test statistic is much more complicated. Therefore, if $p_{\mathrm{naive}}$ is used for decision-making, the false detection error cannot be properly controlled.